Newton is 8 mi east of Oldtown and Littleton is 10 mi northwest of Oldtown. How far is Newton from Littleton?
Answer by Cromlix(4381)

Hi there,
The angle between Newton and
Littleton = 135 degrees
Using Cosine Rule:
Triangle OLN
O^2 = L^2 + N^2 - 2LN Cos 135
O^2 = 10^2 + 8^2 - (2x10x8xCos(135))
O^2 = 277.14
O = 16.65 miles (2 decimal places)
This is the distance between Newton and Littleton.
Hope this helps :-)